Summary

Several camera problems (rotator, temperature ...)were meet. Two hours observation time lost. The seeing is 0.8-1.0" for red channel and 1-1.6" for the blue channel. Focus is smooth and does not take much time. Finally, we took 3 obs for EGS field, 1 ob for Bootes, 2 obs for Herc field

Details

Observers: Xiaohui Fan, Linhua Jiang, Fuyan Bian

Supporting Astronomer: Olga Kuhn

Operator: Aaron Ceranski

All Times UT

First night of the AZ queue.

0130 Sky looks good, a little cloud around horizon, tonight is not photometry night

0410 after a large number of problems with camera, in particular a data probe connection and rotator, which took the last 2.5 hours to fix (we hope), we are finally on the first science field.

0412 surprisingly, the very first lbcpfic took only two iterations to converge; estimated seeing 0.9". starting EGS_4b; the first Y band image is 1" and the first U band is 1.5", the seeing in the rest images is similar to the first one. The last image is bad.

0441 do focus; estimated seeing is 1.2"

0443 starting EGS_4a; the seeing is 1.1" for Y band and 1.6" for U band. The seeing is becoming better. The seeing for the last one is 0.9" for Y band and 1.3 for U band.

0514 do focus; estimated seeing is 0.8

0517 going back EGS_4b; Seeing is 0.9 for Y band and U band (lbcb.20080529.051617.fits) failed to find the guide star, so it is a little trail. The technique image file is lbcbtec.20080529.051643_1.fits, and there is a bright star at the edge (about 10 pixels). Seeing does not change much.

0543 Moving to bootes field : b46r. Do focus; estimated seeing is 0.8". The seeing for the first image is 0.9; the star in the central image looks good, but the stars at the edge are not round. The seeing for the second image is 0.9 and the seeing seems to become worse from chip1 to chip3. The third one ( lbcr.20080529.055719.fits) is bad (the telescope jumped). The seeing becomes worse, 1.0" for the forth one and 1.1 for the fifth one.

0614 do focus; estimated seeing is 0.8"

0617 Starting b01r, The seeing of the first image is 0.9". The rotator does not work again, when taking second image.

0714 Do focus; Crash again

0825 Now the LBC is OK. Going to b01r. Do focus; estimated seeing is 1". The seeing for the first image is 1.1". The red camera stops when taking the second image (Temperature is too high) .

0909 Now LBC is OK. Going to Herc_1_Br. Do focus; estimated seeing is 0.7". The seeing of the first B band image is 1.3" and the first r band image is 1". The shape of the star in the chip 3 is not good. The shape of the star in the second B band image is very bad (lbcb.20080529.092532.fits). The third one is a little better. The fourth one looks good. Seeing becomes worse. The seeing of r band is 1.3 and of B is 1.5. The fifth one looks bad. The sixth one looks OK.

0959 We re do the second, third, fourth one of Herc_1_Br in B band, the seeing is 1.3".

1020 Do focus, estimated seeing is 0.7. Going to Herc_2_Br. 0.8-1'' for r band images and 1.0-1.2" for B band. The shape of the star looks good.

1101 Taking 25 bias images.

1117 Taking twilight flats. skyFlat_B, skyFlat_r, skyflat_U, skyflat_Y When taking the flats, several warnings are encountered, we just ignore them, and the LBC runs smoothly.

-- XiaohuiFan - 29 May 2008
Topic revision: r6 - 30 May 2008, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback